# LLM-Oriented Information Retrieval: A Denoising-First Perspective

### 01 May 2026 v1

@@ 1. Introduction
Another critical challenge is the limited attention and context capacity of current LLMs. Even with the latest models featuring context windows extended to millions of tokens, LLMs struggle to fully utilize very large contexts, especially in reasoning tasks (Liu et al., 2024a; Chen et al., 2024b). Empirical results confirm that beyond a certain point, adding more retrieved passages, especially irrelevant and redundant ones, yields diminishing or even negative returns. Furthermore, noisy or malicious content in the context can actively undermine the model’s reliability (Dai et al., ). A well-known vulnerability is prompt injection, where adversarially crafted text in the retrieved context hijacks the model’s behavior or injects false information (Liu et al., 2023). This highlights that irrelevant or harmful information in the context window isn’t just a waste of computation resources – it can directly degrade output quality or even pose serious risks.
@@ 2. Challenge shifts in the history of IR
First, LLMs have turned retrieval from a destination into an input channel. In retrieval-augmented generation, the consumer is no longer a human browsing for links, but a model requiring actionable evidence to drive generation (Lewis et al., 2020; Wang et al., 2024). Retrieval success is thus redefined by downstream utility: a document is only useful if it positively influences the model’s reasoning trace (Salemi and Zamani, 2024; Dai et al., ).
@@ 3. Denoising in LLM-Oriented Retrieval: A Method Taxonomy
Context construction noise. Noise is often amplified in the last mile between retrieval and generation, when candidates are chunked, merged, and ordered into a single prompt. First, structural and positional noise stems from the mismatch between retrieval ranking and LLM attention; naive concatenation often buries critical evidence in the “lost-in-the-middle” zone, rendering it effectively invisible to the model (Liu et al., 2024a; Wang et al., 2025d). Second, evidence incoherence arises when the retriever aggregates conflicting or redundant passages; this incoherence increases parametric uncertainty and triggers hallucinations, requiring diversity-aware compression rather than simple concatenation (Xie et al., 2023; Wang et al., 2025b). Third, the retrieval channel introduces adversarial noise via indirect prompt injections, where malicious instructions embedded in untrusted documents hijack the generation logic, breaching the system’s safety alignment (Greshake et al., 2023). Consequently, the IR system must treat evidence assembly as an active denoising stage—responsible for curating, ordering, and sanitizing the prompt—rather than a passive data hand-off.
@@
Controlled Indexing3.1)Robust Retrieval3.2)Context Assembly3.3)RetrievalVerification3.4)Closed-Loop Training3.5)Provenance & TrustQuality CanonicalizationTemporal ValidityStructure-aware Indexing Watermark (Kirchenbauer et al., 2023); SynthID (Dathathri et al., 2024); DetectGPT (Mitchell et al., 2023); MAGE (Li et al., 2024b); Spiral (Chen et al., 2024c); Cocktail (Dai et al., 2024b); GLTR (Gehrmann et al., 2019); Model collapse (Shumailov et al., 2024); LLM adoption (Liang et al., 2025a); CopyrightDetective (Zhang et al., 2026a, 2025b) MinHash (Broder, 1997); Dedup data (Lee et al., 2022); CCNet (Wenzek et al., $20$); RefinedWeb (Penedo et al., 2023); FineWeb (Penedo et al., 2024); Dolma (Soldaini et al., 2024) HOH (Ouyang et al., 2025); VersionRAG (Huwiler et al., 2025); EraRAG (Zhang et al., 2025a); Theanine (Ong et al., 2025); STAR-RAG (Zhu et al., 2025c); T-GRAG (Li et al., 2025a) GraphRAG (Edge et al., 2024); GRAG (Hu et al., 2024); LightRAG (Guo et al., 2024); KG2RAG (Zhu et al., 2025b); G-Retriever (He et al., 2024); EDC (Zhang and Soh, 2024); RAKG (Zhang et al., 2025c); RAPTOR (Sarthi et al., 2024); PageIndex (Zhang et al., 2025e); HippoRAG (Gutiérrez et al., 2024) Query TransformationHybrid RetrievalDistractor-aware TrainingFine-grained Matching HyDE (Gao et al., 2023a); Query2Doc (Wang et al., 2023b); Step-Back (Zheng et al., ); GenRead (Yu et al., ); Self-Ask (Press et al., 2023); Enhance-QR (Ye et al., 2023); QR-RAG (Ma et al., 2023a); RaFe (Mao et al., 2024); DMQR-RAG (Li et al., 2024c); Converser (Huang et al., 2023); Decomposition (Mao et al., 2023); MMLF (Kuo et al., 2025); RQ-RAG (Chan et al., ); RL rewrite (Peng et al., 2024; Jiang et al., 2025) Hybrid retrieval (Bruch et al., 2023); SPLADE (Formal et al., 2021); BGE-M3 (Chen et al., 2024a); OOD Hybrid (Chen et al., 2022); Adaptive routing (Mallen et al., 2023); Unsupervised Dense IR (Izacard et al., ) ANCE (Xiong et al., ); RocketQA (Qu et al., 2021); SimLM (Wang et al., 2023a); AR2 (Zhang et al., ); DRAGON (Lin et al., 2023a); INSTRUCTOR (Su et al., 2023); E5 (Wang et al., 2022); RA-DIT (Lin et al., 2023b); Inst. tuning (Wang et al., 2025e; Dai et al., 2024a); RbFT (Tu et al., 2025) ColBERT (Khattab and Zaharia, 2020); ColBERTv2 (Santhanam et al., 2022); XTR (Lee et al., 2023); MUVERA (Dhulipala et al., 2024); ColPali (Faysse et al., ) Selection & PruningContext CompressionLayout OptimizationConflict Resolution Pairwise/Setwise/Listwise reranking (Qin et al., 2024; Zhuang et al., 2024; Reddy et al., 2024); MMR (Carbonell and Goldstein, 1998); DPP (Sarlak and Razi, 2026); Speculative selection (Wang et al., 2025f); FILCO (Wang et al., 2023c); Provence (Chirkova et al., ); DSLR (Hwang et al., 2024); Adaptive-RAG (Jeong et al., 2024) LLMLingua (Jiang et al., 2023a); LongLLMLingua (Jiang et al., 2024); LLMLingua-2 (Pan et al., 2024); RECOMP (Xu et al., 2024); Compression for RAG (Li et al., 2023); Gist Tokens (Mu et al., 2023); ICAE (Ge et al., ); xRAG (Cheng et al., 2024); Token-efficiency objective (Xu et al., 2025) Lost in the Middle (Liu et al., 2024a); Topology-aware ordering (Peysakhovich and Lerer, 2023); OP-RAG (Yu et al., 2024); Irrelevant-context robustness (Shi et al., 2023) Conflict-aware assembly (Xie et al., 2023; Wang et al., 2025b); Chain-of-Note (Yu et al., 2023); Astute RAG (Wang et al., 2025a); Conflict diagnostics (Chen et al., 2024b; Liu et al., 2019); OpenDecoder (Mo et al., 2026) Fine-grained AttributionSafety EvaluationUtility Evaluation FactScore (Min et al., 2023); ALCE (Gao et al., 2023b); GopherCite (Menick et al., 2022); Self-RAG (Asai et al., ); CoVe (Dhuliawala et al., 2024); AlignScore (Zha et al., 2023); MiniCheck (Tang et al., 2024) Prompt-injection defense (Liu et al., 2024c); StruQ (Chen et al., 2025); RobustRAG (Xiang et al., 2024); CAR defense (Weller et al., 2024) RAGChecker (Ru et al., 2024); RAGAS (ES et al., 2023); SePer (Dai et al., ); RAGTruth (Niu et al., 2024); FaithEval (Ming et al., ); SafeRAG (Liang et al., 2025b); Robustness stress test (Chen et al., 2024b) Loop OrchestrationEnd-to-end Policy LearningSelf-evolution IRCoT (Trivedi et al., 2023); ChainRAG (Zhu et al., 2025a); LevelRAG (Zhang et al., 2025f); MindSearch (Chen et al., ); FLARE (Jiang et al., 2023b); DRAGIN (Su et al., 2024); ReAct (Yao et al., 2022); WebGPT (Nakano et al., 2021); TIR (Su et al., 2026) OpenRAG (Zhou and Chen, 2025); RouteRAG (Guo et al., 2025); Search-R1 (searchr1); Toolformer (Schick et al., 2023) Reflexion (Shinn et al., 2023); MemGPT (Packer et al., 2023); MemRL (Zhang et al., 2026b); MemoryBank (Zhong et al., 2024); HippoRAG (Gutiérrez et al., 2024); AutoRAG (Kim et al., 2024); DSPy (Khattab et al., 2024)
@@ 3.2. Robust Retrieval
(3) Distractor-Aware Retriever Training. Enhancing the robustness of dense representations is critical for eliminating noise. Several axes strengthen the retriever training against distractors. Hard negative mining force the encoder to separate genuine evidence from topical near-misses (Qu et al., 2021; Wang et al., 2023a; Lin et al., 2023a), with adversarial frameworks dynamically escalating training difficulty, exposing the retriever to ever-subtler distractor (Zhang et al., ). Secondly, instruction tuning conditions the same embedding space on different task intents, so semantically similar but task-irrelevant passages are systematically downranked (Wang et al., 2025e; Dai et al., 2024a; Su et al., 2023). Finally, RAG-specific objectives further distill LLM-judged utility or defects back into the retriever, aligning what is retrieved with what the generator should actually exploit (Lin et al., 2023b; Tu et al., 2025).
@@ 3.3. Context Assembly
(1) LLM-aware Selection and Pruning. After first-stage retrieval, reranking and pruning act as the first noise gate of context assembly, rigorously filtering irrelevant candidates before they enter the prompt. LLM-based reranking enhances pointwise cross-encoders (Nogueira and Cho, 2019) with holistic, set-level judgments: listwise permutation generation (Sun et al., 2023; Ma et al., 2023b), pairwise comparison (Qin et al., 2024), setwise most-relevant selection (Zhuang et al., 2024), and efficiency-oriented single-token decoding (Reddy et al., 2024). Diversity-aware subset selection moves beyond top-$k$ truncation to balance relevance against redundancy via MMR (Carbonell and Goldstein, 1998) or Determinantal Point Processes (Sarlak and Razi, 2026). Fine-grained context filtering operates at the sentence or token level to surgically remove non-supporting content. Speculative selection instructs the LLM itself to prune irrelevant evidence before reasoning (Wang et al., 2025f); trained pruners (Wang et al., 2023c; Chirkova et al., ) distill comparable signals into lightweight classifiers, and applies sentence-level reranking followed by LLM-based reconstruction of the surviving fragments (Hwang et al., 2024) . Beyond rigid top-$k$ paradigm, adaptive strategies can dynamically size the context based on query complexity (Jeong et al., 2024) or abstain when evidence is insufficient (Penha and Hauff, 2021; Meng et al., 2023; Amiraz et al., 2025).
@@
(2) Long Context Compression. To mitigate noise from verbose or irrelevant tokens, compression techniques perform soft denoising. Extractive compression methods such as the LLMLingua family (Jiang et al., 2023a, 2024; Pan et al., 2024) use small models to score token informativeness and drop low-signal tokens. Abstractive compression synthesizes documents into concise meta-summaries or task-specific representations (Xu et al., 2024; Li et al., 2023). A complementary line, soft-prompt compression, encodes retrieved passages into continuous embedding vectors—such as Gist Tokens (Mu et al., 2023), ICAE (Ge et al., ), and xRAG (Cheng et al., 2024)—achieving extreme compression while preserving task performance. All three paradigms aim to maximize information gain per token (Xu et al., 2025), increasing the logical capacity of the context window.
@@
(3) Layout Optimization to Mitigate Position Bias. LLMs exhibit non-uniform attention distributions, often favoring the beginning and end of the context while neglecting the middle—the “lost in the middle” phenomenon (Liu et al., 2024a). Denoising in this dimension involves topology-aware arrangement, where retrieved passages are reordered to place the most critical evidence at attention peaks (Peysakhovich and Lerer, 2023). Rather than a naive concatenation based on retrieval scores, optimal layout strategies dynamically structure the prompt to align with the model’s inductive biases, ensuring that high-signal evidence is not statistically drowned out by surrounding context (Shi et al., 2023).
@@
(4) Knowledge Conflict Resolution. Retrieved evidence is often internally inconsistent due to temporal drift, source disagreement, or duplicated paraphrases; naively flattening such passages into a single prompt can lead to hallucinations or reasoning paralysis (Chen et al., 2024b; Liu et al., 2019). Conflict-aware assembly therefore treats contradiction as a first-class denoising target: competing claims can be detected and clustered into explicit alternatives, or routed into a structured prompt that forces the model to adjudicate discrepancies rather than averaging them away (Xie et al., 2023; Wang et al., 2025b). Complementarily, evidence grading methods such as Chain-of-Note (Yu et al., 2023) generate passage-level reading notes and labels (e.g., supporting vs. irrelevant vs. contradictory), enabling the assembler to prune low-signal passages and to expose disagreements with clear provenance before final synthesis. OpenDecoder(Mo et al., 2026) incorporates retrieval relevance signals in answer decoding, modulating retrieval and generation with an internal attention network.

### 17 May 2026 v2

@@ 1. Introduction
Another critical challenge is the limited attention and context capacity of current LLMs. Even with the latest models featuring context windows extended to millions of tokens, LLMs struggle to fully utilize very large contexts, especially in reasoning tasks (Liu et al., 2024a; Chen et al., 2024b). Empirical results confirm that beyond a certain point, adding more retrieved passages, especially irrelevant and redundant ones, yields diminishing or even negative returns. Furthermore, noisy or malicious content in the context can actively undermine the model’s reliability (Dai et al., 2025). A well-known vulnerability is prompt injection, where adversarially crafted text in the retrieved context hijacks the model’s behavior or injects false information (Liu et al., 2023). This highlights that irrelevant or harmful information in the context window isn’t just a waste of computation resources – it can directly degrade output quality or even pose serious risks.
@@ 2. Challenge shifts in the history of IR
First, LLMs have turned retrieval from a destination into an input channel. In retrieval-augmented generation, the consumer is no longer a human browsing for links, but a model requiring actionable evidence to drive generation (Lewis et al., 2020; Wang et al., 2024a). Retrieval success is thus redefined by downstream utility: a document is only useful if it positively influences the model’s reasoning trace (Salemi and Zamani, 2024; Dai et al., 2025).
@@ 3. Denoising in LLM-Oriented Retrieval: A Method Taxonomy
Context construction noise. Noise is often amplified in the last mile between retrieval and generation, when candidates are chunked, merged, and ordered into a single prompt. First, structural and positional noise stems from the mismatch between retrieval ranking and LLM attention; naive concatenation often buries critical evidence in the “lost-in-the-middle” zone, rendering it effectively invisible to the model (Liu et al., 2024a; Wang et al., 2025e). Second, evidence incoherence arises when the retriever aggregates conflicting or redundant passages; this incoherence increases parametric uncertainty and triggers hallucinations, requiring diversity-aware compression rather than simple concatenation (Xie et al., 2023; Wang et al., 2025c). Third, the retrieval channel introduces adversarial noise via indirect prompt injections, where malicious instructions embedded in untrusted documents hijack the generation logic, breaching the system’s safety alignment (Greshake et al., 2023). Consequently, the IR system must treat evidence assembly as an active denoising stage—responsible for curating, ordering, and sanitizing the prompt—rather than a passive data hand-off.
@@
Controlled Indexing3.1)Robust Retrieval3.2)Context Assembly3.3)RetrievalVerification3.4)Closed-Loop Training3.5)Provenance & TrustQuality CanonicalizationTemporal ValidityStructure-aware Indexing Watermark (Kirchenbauer et al., 2023); SynthID (Dathathri et al., 2024); DetectGPT (Mitchell et al., 2023); MAGE (Li et al., 2024b); Spiral (Chen et al., 2024c); Cocktail (Dai et al., 2024b); GLTR (Gehrmann et al., 2019); Model collapse (Shumailov et al., 2024); LLM adoption (Liang et al., 2025a); CopyrightDetective (Zhang et al., 2026a, 2025b) MinHash (Broder, 1997); Dedup data (Lee et al., 2022); CCNet (Wenzek et al., $20$); RefinedWeb (Penedo et al., 2023); FineWeb (Penedo et al., 2024); Dolma (Soldaini et al., 2024) HOH (Ouyang et al., 2025); VersionRAG (Huwiler et al., 2025); EraRAG (Zhang et al., 2025a); Theanine (Ong et al., 2025); STAR-RAG (Zhu et al., 2025c); T-GRAG (Li et al., 2025a) GraphRAG (Edge et al., 2024); GRAG (Hu et al., 2024); LightRAG (Guo et al., 2024); KG2RAG (Zhu et al., 2025b); G-Retriever (He et al., 2024); EDC (Zhang and Soh, 2024); RAKG (Zhang et al., 2025c); RAPTOR (Sarthi et al., 2024); PageIndex (Zhang et al., 2025e); HippoRAG (Gutiérrez et al., 2024) Query TransformationHybrid RetrievalDistractor-aware TrainingFine-grained Matching HyDE (Gao et al., 2023a); Query2Doc (Wang et al., 2023b); Step-Back (Zheng et al., ); GenRead (Yu et al., ); Self-Ask (Press et al., 2023); Enhance-QR (Ye et al., 2023); QR-RAG (Ma et al., 2023a); RaFe (Mao et al., 2024); DMQR-RAG (Li et al., 2024c); Converser (Huang et al., 2023); Decomposition (Mao et al., 2023); MMLF (Kuo et al., 2025); RQ-RAG (Chan et al., ); RL rewrite (Peng et al., 2024; Jiang et al., 2025) Hybrid retrieval (Bruch et al., 2023); SPLADE (Formal et al., 2021); BGE-M3 (Chen et al., 2024a); OOD Hybrid (Chen et al., 2022); Adaptive routing (Mallen et al., 2023); Unsupervised Dense IR (Izacard et al., ) ANCE (Xiong et al., ); RocketQA (Qu et al., 2021); SimLM (Wang et al., 2023a); AR2 (Zhang et al., ); DRAGON (Lin et al., 2023a); INSTRUCTOR (Su et al., 2023); E5 (Wang et al., 2022); RA-DIT (Lin et al., 2023b); Inst. tuning (Wang et al., 2025f; Dai et al., 2024a); RbFT (Tu et al., 2025) ColBERT (Khattab and Zaharia, 2020); ColBERTv2 (Santhanam et al., 2022); XTR (Lee et al., 2023); MUVERA (Dhulipala et al., 2024); ColPali (Faysse et al., ) Selection & PruningContext CompressionLayout OptimizationConflict Resolution Pairwise/Setwise/Listwise reranking (Qin et al., 2024; Zhuang et al., 2024; Reddy et al., 2024); MMR (Carbonell and Goldstein, 1998); DPP (Sarlak and Razi, 2026); Speculative selection (Wang et al., 2025g); FILCO (Wang et al., 2023c); Provence (Chirkova et al., ); DSLR (Hwang et al., 2024); Adaptive-RAG (Jeong et al., 2024) LLMLingua (Jiang et al., 2023a); LongLLMLingua (Jiang et al., 2024); LLMLingua-2 (Pan et al., 2024); RECOMP (Xu et al., 2024a); Compression for RAG (Li et al., 2023); Gist Tokens (Mu et al., 2023); ICAE (Ge et al., ); xRAG (Cheng et al., 2024); Token-efficiency objective (Xu et al., 2025) Lost in the Middle (Liu et al., 2024a; Tian et al., 2025); Mechanistic analysis (Yu et al., 2025b; Wang et al., 2024b); Positional calibration (Hsieh et al., 2024); Topology-aware reordering (Peysakhovich and Lerer, 2023; Byerly and Khashabi, 2026); OP-RAG (Yu et al., 2024); Irrelevant-context robustness (Shi et al., 2023) conflict-aware prompting (Xie et al., 2023; Wang et al., 2025c), Chain-of-Note (Yu et al., 2023), Astute RAG (Wang et al., 2025a); AdaCAD (Shi et al., 2024; Wang et al., 2025b), OpenDecoder (Mo et al., 2026); RGB (Chen et al., 2024b); ConflictBank (Su et al., 2024b); WikiContradict (Hou et al., 2024) Fine-grained AttributionSafety EvaluationUtility Evaluation FactScore (Min et al., 2023); ALCE (Gao et al., 2023b); GopherCite (Menick et al., 2022); CoVe (Dhuliawala et al., 2024); AlignScore (Zha et al., 2023); MiniCheck (Tang et al., 2024) Prompt-injection defense (Liu et al., 2024c); StruQ (Chen et al., 2025); RobustRAG (Xiang et al., 2024); CAR defense (Weller et al., 2024) RAGChecker (Ru et al., 2024); RAGAS (ES et al., 2023); SePer (Dai et al., 2025); RAGTruth (Niu et al., 2024); FaithEval (Ming et al., ); SafeRAG (Liang et al., 2025b); Robustness stress test (Chen et al., 2024b) Loop OrchestrationEnd-to-end Policy LearningSelf-evolution IRCoT (Trivedi et al., 2023); ChainRAG (Zhu et al., 2025a); LevelRAG (Zhang et al., 2025f); MindSearch (Chen et al., ); FLARE (Jiang et al., 2023b); DRAGIN (Su et al., 2024a); ReAct (Yao et al., 2022); WebGPT (Nakano et al., 2021); TIR (Su et al., 2026) OpenRAG (Zhou and Chen, 2025); RouteRAG (Guo et al., 2025); Search-R1 (searchr1); Toolformer (Schick et al., 2023) Reflexion (Shinn et al., 2023); MemGPT (Packer et al., 2023); MemRL (Zhang et al., 2026b); MemoryBank (Zhong et al., 2024); HippoRAG (Gutiérrez et al., 2024); AutoRAG (Kim et al., 2024); DSPy (Khattab et al., 2024); Self-RAG (Asai et al., )
@@ 3.2. Robust Retrieval
(3) Distractor-Aware Retriever Training. Enhancing the robustness of dense representations is critical for eliminating noise. Several axes strengthen the retriever training against distractors. Hard negative mining force the encoder to separate genuine evidence from topical near-misses (Qu et al., 2021; Wang et al., 2023a; Lin et al., 2023a), with adversarial frameworks dynamically escalating training difficulty, exposing the retriever to ever-subtler distractor (Zhang et al., ). Secondly, instruction tuning conditions the same embedding space on different task intents, so semantically similar but task-irrelevant passages are systematically downranked (Wang et al., 2025f; Dai et al., 2024a; Su et al., 2023). Finally, RAG-specific objectives further distill LLM-judged utility or defects back into the retriever, aligning what is retrieved with what the generator should actually exploit (Lin et al., 2023b; Tu et al., 2025).
@@ 3.3. Context Assembly
(1) LLM-aware Selection and Pruning. After first-stage retrieval, reranking and pruning act as the first noise gate of context assembly, rigorously filtering irrelevant candidates before they enter the prompt. LLM-based reranking enhances pointwise cross-encoders (Nogueira and Cho, 2019) with holistic, set-level judgments: listwise permutation generation (Sun et al., 2023; Ma et al., 2023b), pairwise comparison (Qin et al., 2024), setwise most-relevant selection (Zhuang et al., 2024), and efficiency-oriented single-token decoding (Reddy et al., 2024). Diversity-aware subset selection moves beyond top-$k$ truncation to balance relevance against redundancy via MMR (Carbonell and Goldstein, 1998) or Determinantal Point Processes (Sarlak and Razi, 2026). Fine-grained context filtering operates at the sentence or token level to surgically remove non-supporting content. Speculative selection instructs the LLM itself to prune irrelevant evidence before reasoning (Wang et al., 2025g); trained pruners (Wang et al., 2023c; Chirkova et al., ) distill comparable signals into lightweight classifiers, and applies sentence-level reranking followed by LLM-based reconstruction of the surviving fragments (Hwang et al., 2024) . Beyond rigid top-$k$ paradigm, adaptive strategies can dynamically size the context based on query complexity (Jeong et al., 2024) or abstain when evidence is insufficient (Penha and Hauff, 2021; Meng et al., 2023; Amiraz et al., 2025).
@@
(2) Long Context Compression. To mitigate noise from verbose or irrelevant tokens, compression techniques perform soft denoising. Extractive compression methods such as the LLMLingua family (Jiang et al., 2023a, 2024; Pan et al., 2024) use small models to score token informativeness and drop low-signal tokens. Abstractive compression synthesizes documents into concise meta-summaries or task-specific representations (Xu et al., 2024a; Li et al., 2023). A complementary line, soft-prompt compression, encodes retrieved passages into continuous embedding vectors—such as Gist Tokens (Mu et al., 2023), ICAE (Ge et al., ), and xRAG (Cheng et al., 2024)—achieving extreme compression while preserving task performance. All three paradigms aim to maximize information gain per token (Xu et al., 2025), increasing the logical capacity of the context window.
@@
(3) Layout Optimization to Mitigate Position Bias. LLMs exhibit non-uniform attention, favoring the beginning and end of the context while neglecting the middle—the “lost in the middle” phenomenon (Liu et al., 2024a)—a bias traced to causal masking and specific hidden dimensions (Yu et al., 2025b; Wang et al., 2024b), and amplified when relevant pieces are spaced far apart (Tian et al., 2025). Denoising involves topology-aware reordering to place critical evidence at attention peaks (Peysakhovich and Lerer, 2023), positional calibration to flatten the U-shaped curve (Hsieh et al., 2024), and inference-time shuffling with aggregation (Byerly and Khashabi, 2026). These strategies dynamically structure the prompt to align with the model’s inductive biases, preventing high-signal evidence from being drowned out (Shi et al., 2023).
@@
(4) Knowledge Conflict Resolution. Retrieved evidence is often internally inconsistent due to temporal drift, source disagreement, or duplicated paraphrases; naively flattening such passages into a single prompt can lead to hallucinations or reasoning paralysis (Chen et al., 2024b; Liu et al., 2019; Xu et al., 2024b), benchmarked by (Hou et al., 2024; Su et al., 2024b). At assembly time, conflict-aware methods detect competing claims and route them into structured prompts that force the model to adjudicate discrepancies rather than averaging them away (Xie et al., 2023; Wang et al., 2025c). Complementarily, evidence grading methods such as Chain-of-Note (Yu et al., 2023) generate passage-level reading notes (supporting vs. irrelevant vs. contradictory), while Astute RAG (Wang et al., 2025a) iteratively consolidates model’s own parametric knowledge with retrieved passages according to source reliability, filtering inconsistencies before generation. During generation, decoding-time interventions offer a second line of defense: contrastive decoding amplifies contextual over parametric signals (Shi et al., 2024), adaptive variants dynamically adjust the contrast per token to avoid over-correction of conflict (Wang et al., 2025b), and OpenDecoder (Mo et al., 2026) injects explicit passage quality indicators into the attention computation to steer generation toward higher-reliability passages.